OpenStudy (anonymous):

what is the solution to the system of equations? -4x+4y-2z=-8 -3x-y+4z=0 2x-2y+3z=-4 possibles are (-3,-7,-4) (1,-7,-2) (3,5,-4) (-3,0,5)

OpenStudy (anonymous):

i can solve this, but after an hour because i still have my class.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(-3,-7,-4) 12-28+8=-8 9-7=2

OpenStudy (anonymous):

(1,-7,-2) -4-14+4=-14

OpenStudy (anonymous):

(3,5,-4) -15+20+8=13

OpenStudy (anonymous):

(-3,0,5) 12+0-10=2

OpenStudy (anonymous):

first we will are going to find equations with two variables, (-3x-y+4z=0)*4 -12x-4y+16z=0 -4x+4y-2z=-8, add this equations. you will get: -16x+14z=-8 (eqn. 4) (-3x-y+4z=0)*-2 6x+2y-8z=0 2x-2y+3z=-4, add you will get: 8x-5z=-4 (eqn. 5) add eqn 4 and 5 -16x+14z=-8 (8x-5z=-4)*2 -----> 16x-10z=-8 you will get 4z=-16 z=-4 -4x+4y-2(-4)=-8 -4x+4y+8=-8 -4x+4y=-8-8 -4x+4y=-16 (eqn. 6) z=-4 -3x-y+4(-4)=0 -3x-y-16=0 -3x-y=16 (eqn. 7) add eqn 6 and 7. -4x+4y=-16 (-3x-y=16)*4-----> -12x-4y=64 -16x=48 x=-3 then, x=-3 and z=-4 -4x+4y-2z=-8 -4(-3)+4y-2(-4)=-8 12+4y+8=-8 4y=-8-8-12 4y=-28 y=-7 So, x=-3, y=-7 and z=-4